Nancy is reviewing her November credit card statement. Her beginning balance was $610 and she made a $150 payment on November 10. She made purchases of $100 on November 5, $120 on Novemeber 15, and $45 on November 30, Her APR is 15% and the interest was charged using the average daily balance method, including current purchases, which considers the day of a charge or credit. Answer parts 1 through 5.

`1. With 30 Days in November, what is the average daily balance

2. Calculate interest in November

3. What was the unpaid balance for November after interest is charged?

4. If Nancy's account instead used the unpaid balance method, calculate the finance charge and the new balance for the month of November.

5. The new balance for the month of November would be
Answer #1

1:

Period Number of days Balance Days* Balance
1 Nov to 4 Nov 4 610 2440
5 Nov to 9 Nov 5 710 3550
10Nov to 14 Nov 5 560 2800
15 Nov to 29 Nov 15 680 10200
30-Nov 1 725 725
TOTAL 30 3285 19715

Average daily balance = Sum of days*Daily balance/ Number of days

= 19715/ 30 = $657.17

2: Interest in Novermber = Daily balance*APR*30/365

= 657.17* 15%*30/365

=$8.10

3: Unpaid balance in November = Balance on 30 Nov + Interest

= 725+ 8.10

= $733.10

4: Unpaid balance Method

Finance charge =

Unpaid balance*APR*30/365

= 725* 15%*30/365 = 8.94

5: New balance = 725+ 8.94 = 733.94